The total area of a rectangle is 120 feet squared. The length of the rectangle is three times the width, plus 9. Using the above information, write out the quadratic equation you would use to solve: what is the length and width of the rectangle? Finally, what is the length and width of the rectangle?

How to solve ?

Answer by jorel1380(3719)

Let w be the width. Then the length would be 3w+9. So:
w(3w+9)=120
3w²+9w-120=0
w²+3w-40=0
(w+8)(w-5)=0
w=5 or -8
The length and width of the rectangle is 24 x 5 ft. ☺☺☺☺